OpenStudy (anonymous):

Part 1 [3 points]: Find the solution to the equation 64(3 – x) = 42x. Part 2 [4 points]: Use complete sentences to explain the method used to solve this equation. This is my last question so I can finish this semester exam. Thank you if anyone helps! :)

OpenStudy (anonymous):

by simplifying you get 192-106x=0 therefore, x = 96/53 or 1.81

OpenStudy (anonymous):

can you elaborate more?

OpenStudy (anonymous):

explain every step that you need to perform to solve this. so basically explain what you did in part 1..

OpenStudy (anonymous):

64^(3 – x) = 4^(2x) 64 = 4^3 4^(3(3 – x)) = 4^2x 3(3 – x) = 2x 9 - 3x = 2x 9 = 5x x = 9/5 I got this

OpenStudy (anonymous):

*thumbs up*

OpenStudy (anonymous):

oh then 9/5 is 1.8 gt it thanks

OpenStudy (anonymous):

\[\LARGE 64(3-x)=42x \] multiply 64 with both numbers into bracket. (with 3 and -x) \[\LARGE 64\times 3+64\times (-x)=42x \] now multiply ... \[\LARGE 192 -64x=42x \] now move -64x to the right side... (and its sign will change into +) \[\LARGE 192 =42x+64x \] now sum 42x and 64x \[\LARGE 192 =106x \] now divide both sides by 106 \[\LARGE \frac{192}{106} =\frac{106x}{106 } \] \[\LARGE 1.81 =x \]

OpenStudy (anonymous):

Kreshnik thank you so much! Helped a lot :D